# MissionGNN: Hierarchical Multimodal GNN-based Weakly Supervised Video Anomaly Recognition with Mission-Specific Knowledge Graph Generation

### v1

@@ IV-D Evaluation on VAD and VAR
Method Class mAUC AUC Abuse Arrest Arson Assault Burglary Explosion Fighting RoadAcc. Robbery Shooting Shoplifting Stealing Vandalism RTFM [38] 79.99 62.57 90.53 82.27 85.53 92.76 85.21 90.31 81.17 82.82 92.56 90.23 87.20 84.86 84.03 S3R [47] 86.38 68.45 92.19 93.55 86.91 93.55 81.69 85.03 82.07 85.32 91.64 94.59 83.82 86.55 85.99 SSRL [17] 95.33 79.26 93.27 91.74 89.06 92.25 87.36 80.24 87.75 84.5 92.31 94.22 88.17 88.88 87.43 ActionCLIP [43] 91.88 90.47 89.21 86.87 81.31 94.08 83.23 94.34 82.82 70.53 91.60 94.06 89.89 87.71 82.30 AnomalyCLIP [53] 75.03 94.56 96.66 94.80 90.08 94.79 88.76 93.30 86.85 87.45 89.47 97.00 89.78 90.66 86.36 CLIP zero-shot [29] 57.37 80.65 91.72 80.83 74.34 90.31 83.54 87.46 70.22 63.99 71.21 45.49 66.45 74.12 58.63 MissionGNN (Ours) 93.68 87.64 94.15 95.86 92.31 94.73 89.44 97.16 81.40 90.57 92.94 97.74 81.77 91.49 84.48
@@
Method Class mAP AP Abuse CarAccident Explosion Fighting Riot Shooting RTFM [38] 9.25 25.36 53.53 61.73 90.38 18.01 43.04 77.81 S3R [47] 2.63 23.82 45.29 49.88 90.41 4.34 36.06 80.26 ActionCLIP [43] 2.73 25.15 55.28 58.09 87.31 12.87 40.24 61.01 AnomalyCLIP [53] 6.10 31.31 68.75 71.44 92.74 26.13 49.41 78.51 CLIP zero-shot [29] 0.32 12.21 22.26 25.25 66.60 1.26 21.32 27.21 MissionGNN (Ours) 10.98 8.91 42.49 32.69 81.64 11.78 31.41 98.42
@@ V Discussion
Trade-off between VAD and VAR performance in different dataset settings. UCF-Crime dataset comprises videos recorded exclusively by fixed surveillance cameras in real-world settings, where the majority of video frames exhibit consistent backgrounds. In contrast, the XD-Violence dataset offers a diverse range of videos, spanning from movies to basketball games, resulting in frequent changes in camera position between frames. This variability poses a challenge to the ability of our proposed method to effectively capture short-term recognition. MissionGNN outperforms state-of-the-art VAR tasks in the UCF-Crime dataset but lags slightly behind in VAD tasks. Conversely, AnomalyCLIP [53] demonstrates superior performance in terms of VAR within the XD-Violence dataset, while our framework significantly outperforms state-of-the-art VAD tasks in the same dataset. AnomalyCLIP’s advantage stems from its comprehensive analysis of all video frames, incorporating information from past to future by processing entire video snippets. However, training with future values can lead to overly optimistic results and may not be pragmatic for real-world deployment.
Inspired by human intuition, MissionGNN is constrained to classify 30 frames. The primary limitation of our proposed method lies in its inability to capture long-term dependencies for anomaly detection. However, in real-world applications such as violence alerting, timely responsiveness is crucial for effective anomaly detection. Therefore, there is a critical need for anomaly detection models with rapid response capabilities. Our proposed method achieves superior results compared to state-of-the-art models while meeting these requirements for both VAD and VAR, making a significant contribution to the field.
@@ Acknowledgements
This work was supported in part by the DARPA Young Faculty Award, the National Science Foundation (NSF) under Grants #2127780, #2319198, #2321840, #2312517, and #2235472, the Semiconductor Research Corporation (SRC), the Office of Naval Research through the Young Investigator Program Award, and Grants #N00014-21-1-2225 and #N00014-22-1-2067. Additionally, support was provided by the Air Force Office of Scientific Research under Award #FA9550-22-1-0253, along with generous gifts from Xilinx and Cisco.

### v2

@@ IV-D Evaluation on VAD and VAR
IV-E Evaluation on Real-time VAD and VAR
A key advantage of the MissionGNN framework is its ability to operate with only 30 consecutive frames during both training and inference, making it well-suited for real-time anomaly detection and recognition. In contrast, AnomalyCLIP, which achieves higher performance in some evaluations, requires the entire video sequence for both training and inference, limiting its practical use in real-time applications such as intelligent surveillance. To illustrate this distinction, we conducted a comparative evaluation on the UCF-Crime and XD-Violence datasets, constraining AnomalyCLIP to the same 30-frame input as our model. The results, shown in Table II and Table III, highlight the impact of this limitation on AnomalyCLIP’s performance. In the VAR task, restricting the frame count caused a significant performance drop across all anomaly classes for AnomalyCLIP, resulting in reductions of 4.5% in mAUC and 14.29% in mAP scores across both datasets. This adjustment altered the performance comparison between the two methods. On the UCF-Crime dataset, MissionGNN exhibited a 6.23% improvement over AnomalyCLIP, further widening the performance gap. On the XD-Violence dataset, however, the gap narrowed by 3.71%, reducing the original 18% difference, which emphasizes the limitations of previous approaches in real-time settings. In the VAD task, MissionGNN outperformed AnomalyCLIP on both datasets, with gains of 1.13% in AUC and 32.91% in AP scores. These findings underscore the performance degradation of AnomalyCLIP when adapted for real-time analysis, as reflected in its lower AUC and AP scores on both datasets, reinforcing the practical advantages of our proposed framework for real-time anomaly detection and recognition.
@@ IV-E Evaluation on Real-time VAD and VAR
Method Class mAUC AUC Abuse Arrest Arson Assault Burglary Explosion Fighting RoadAcc. Robbery Shooting Shoplifting Stealing Vandalism CLIP zero-shot [22] 57.37 80.65 91.72 80.83 74.34 90.31 83.54 87.46 70.22 63.99 71.21 45.49 66.45 74.12 58.63 RTFM [29] 79.99 62.57 90.53 82.27 85.53 92.76 85.21 90.31 81.17 82.82 92.56 90.23 87.20 84.86 84.03 S3R [36] 86.38 68.45 92.19 93.55 86.91 93.55 81.69 85.03 82.07 85.32 91.64 94.59 83.82 86.55 85.99 SSRL [13] 95.33 79.26 93.27 91.74 89.06 92.25 87.36 80.24 87.75 84.5 92.31 94.22 88.17 88.88 87.43 ActionCLIP [34] 91.88 90.47 89.21 86.87 81.31 94.08 83.23 94.34 82.82 70.53 91.60 94.06 89.89 87.71 82.30 AnomalyCLIP [42] 75.03 94.56 96.66 94.80 90.08 94.79 88.76 93.30 86.85 87.45 89.47 97.00 89.78 90.66 86.36 AnomalyCLIP [42] w/ Frame Restriction 69.51 (-5.52) 90.03 (-4.53) 92.44 (-4.22) 88.47 (-6.33) 82.38 (-7.7) 90.44 (-4.35) 85.83 (-2.93) 90.64 (-2.66) 82.22 (-4.63) 83.21 (-4.24) 87.43 (-2.04) 93.95 (-3.05) 83.48 (-6.3) 86.15 (-4.5) 83.35 (-3.01) MissionGNN (Ours) 93.68 92.95 94.15 96.89 92.31 94.79 92.38 97.16 81.40 90.57 92.94 97.74 84.03 92.38 84.48
@@
Method Class mAP AP Abuse CarAccident Explosion Fighting Riot Shooting CLIP zero-shot [22] 0.32 12.21 22.26 25.25 66.60 1.26 21.32 27.21 RTFM [29] 9.25 25.36 53.53 61.73 90.38 18.01 43.04 77.81 S3R [36] 2.63 23.82 45.29 49.88 90.41 4.34 36.06 80.26 ActionCLIP [34] 2.73 25.15 55.28 58.09 87.31 12.87 40.24 61.01 AnomalyCLIP [42] 6.10 31.31 68.75 71.44 92.74 26.13 49.41 78.51 AnomalyCLIP [42] w/ Frame Restriction 4.2 (-1.9) 21.24 (-10.07) 48.34 (-20.41) 47.08 (-24.36) 73.33 (-19.41) 16.53 (-9.6) 35.12 (-14.29) 65.51 (-13) MissionGNN (Ours) 10.98 8.91 42.49 34.06 81.64 15.03 32.18 98.42
@@ V Discussion
Significant Performance Drop of the Previous State-of-the-Art Model on XD-Violence in Real-Time Scenarios. The notable performance drop of AnomalyCLIP on the XD-Violence dataset in real-time scenarios stems from the dataset’s highly diverse and dynamic nature. XD-Violence features a wide range of video types, including fast-paced scenes from movies and sports events, where frequent changes in camera angles and scene transitions are common. These scene swaps make it unrealistic for real-time applications, as relevant frames may be separated by long periods within the video sequence. AnomalyCLIP performs well in this dataset because it processes entire video sequences, allowing it to consider both past and future frames, effectively capturing long-term relationships and subtle anomalies spread across disparate time segments. This capability is particularly advantageous in offline settings, where the model can access all frames at once, contributing to its superior performance in VAR tasks. However, in real-time applications, where future frames are not available, this reliance on long-term temporal context becomes a major limitation. When AnomalyCLIP is constrained to operate within a 30-frame window, as required for real-time analysis, it experiences a significant performance drop, since it can no longer rely on distant future or past frames to refine its predictions. In contrast, MissionGNN is designed specifically for real-time scenarios, where timely and efficient anomaly detection is crucial. By focusing on short-term relationships within the immediate frame window, MissionGNN is more suitable for real-world applications like violence alerting, where multiple scenes and unpredictable events make it impractical to rely on future information. Although MissionGNN may not capture long-term dependencies as effectively, its real-time responsiveness makes it more practical in scenarios requiring immediate detection. This comparison highlights the trade-off between models like AnomalyCLIP, optimized for comprehensive analysis with long-term context, and those like MissionGNN , which are tailored for efficient, real-time anomaly detection despite scene-swapping challenges in datasets like XD-Violence.
@@ Acknowledgements
This work was supported in part by the DARPA Young Faculty Award, the National Science Foundation (NSF) under Grants #2127780, #2319198, #2321840, #2312517, and #2235472, the Semiconductor Research Corporation (SRC), the Office of Naval Research through the Young Investigator Program Award, and Grants #N00014-21-1-2225 and #N00014-22-1-2067, and Army Research Office Grant #W911NF2410360. Additionally, support was provided by the Air Force Office of Scientific Research under Award #FA9550-22-1-0253, along with generous gifts from Xilinx and Cisco.